# Towards Fully Exploiting LLM Internal States to Enhance Knowledge Boundary Perception

### v1

@@ Abstract
Large language models (LLMs) exhibit impressive performance across diverse tasks but often struggle to accurately gauge their knowledge boundaries, leading to confident yet incorrect responses. This paper explores leveraging LLMs’ internal states to enhance their perception of knowledge boundaries from efficiency and risk perspectives. We investigate whether LLMs can estimate their confidence using internal states before response generation, potentially saving computational resources. Our experiments on datasets like Natural Questions, HotpotQA, and MMLU reveal that LLMs demonstrate significant pre-generation perception, which is further refined post-generation, with perception gaps remaining stable across varying conditions. To mitigate risks in critical domains, we introduce Consistency-based Confidence Calibration ($C^{3}$), which assesses confidence consistency through question reformulation. $C^{3}$ significantly improves LLMs’ ability to recognize their knowledge gaps, enhancing the unknown perception rate by 5.6% on NQ and 4.9% on HotpotQA. Our findings suggest that pre-generation confidence estimation can optimize efficiency, while $C^{3}$ effectively controls output risks, advancing the reliability of LLMs in practical applications [^1] The code can be found at GitHub Repository..
@@ 1 Introduction
To enhance LLMs’ perception of what they do not know, we introduce $C^{3}$ (Consistency-based Confidence Calibration), inspired by human behavior, where repeated probing is used to detect inconsistencies and potential deception. $C^{3}$ leverages confidence consistency: if a model is confident in answering a question but loses confidence when the question format changes, this inconsistency signals potential uncertainty, indicating that the model may be overconfident.
@@ 4.1 Experimental Setup
To investigate the factors influencing the difference in the model’s perception before and after generation, we analyze the effects of question difficulty, question format, and the amount of training data. The specific settings are detailed in Paragraph Datasets 4.1. Additionally, to examine whether the information content in generated responses impacts the gap in LLMs’ perception, we employ two strategies: Vanilla, where the model is simply asked to provide the correct answer, and Chain-of-Thought Wei et al. (2022) (COT), where the model first outputs its reasoning process before providing the final answer.
@@ 4.2 Results and Analysis
5 $C^{3}$: Consistency-Based Confidence Calibration
@@ 5.1 Overview
If someone truly knows the correct answer, they will remain confident in their ability to answer the question correctly, even when the question is asked in different ways. Inspired by this, we think a model may be overconfident if the model is confident in its ability to answer a question correctly but loses confidence when the question format changes. In such cases, the model’s original confidence should be calibrated. This approach, which reduces overconfidence by leveraging the consistency of the model’s confidence across differently phrased questions, is termed $C^{3}$ (Consistency-based Confidence Calibration). We focus on calibrating the model’s confidence on free-form questions cause they are the most commonly used question format in practice.
@@ 6 Conclusion
In this paper, we first examine LLMs’ ability to assess the factuality of their responses using internal states before and after response generation. Our findings show that LLMs can predict the correctness of their answers prior to generation, providing a cost-efficient approach that avoids inference, with this ability further enhanced post-generation. Next, we introduce $C^{3}$ (Consistency-based Confidence Calibration), a method to refine the model’s perception after response generation. Experimental results demonstrate that $C^{3}$ substantially improves LLMs’ ability to recognize what they do not know and consistently enhances their overall perception.

### v2

@@ Abstract
Large language models (LLMs) exhibit impressive performance across diverse tasks but often struggle to accurately gauge their knowledge boundaries, leading to confident yet incorrect responses. This paper explores leveraging LLMs’ internal states to enhance their perception of knowledge boundaries from efficiency and risk perspectives. We investigate whether LLMs can estimate their confidence using internal states before response generation, potentially saving computational resources. Our experiments on datasets like Natural Questions, HotpotQA, and MMLU reveal that LLMs demonstrate significant pre-generation perception, which is further refined post-generation, with perception gaps remaining stable across varying conditions. To mitigate risks in critical domains, we introduce Confidence Consistency-based Calibration ($C^{3}$), which assesses confidence consistency through question reformulation. $C^{3}$ significantly improves LLMs’ ability to recognize their knowledge gaps, enhancing the unknown perception rate by 5.6% on NQ and 4.9% on HotpotQA. Our findings suggest that pre-generation confidence estimation can optimize efficiency, while $C^{3}$ effectively controls output risks, advancing the reliability of LLMs in practical applications [^1] The code can be found at GitHub Repository..
@@ 1 Introduction
To enhance LLMs’ perception of what they do not know, we introduce $C^{3}$ (Confidence Consistency-based Calibration), inspired by human behavior, where repeated probing is used to detect inconsistencies and potential deception. $C^{3}$ leverages confidence consistency: if a model is confident in answering a question but loses confidence when the question format changes, this inconsistency signals potential uncertainty, indicating that the model may be overconfident.
@@ 4.1 Experimental Setup
We hypothesize that question difficulty, question format, and the amount of training data may influence the gap between the model’s pre- and post-generation perception. Therefore, we investigate along these dimensions. The specific settings are detailed in Paragraph Datasets 4.1. Additionally, to examine whether the information content in generated responses impacts the gap in LLMs’ perception, we employ two strategies: Vanilla, where the model is simply asked to provide the correct answer, and Chain-of-Thought Wei et al. (2022) (COT), where the model first outputs its reasoning process before providing the final answer.
@@ 4.2 Results and Analysis
5 $C^{3}$: Confidence Consistency-Based Calibration
@@ 5.1 Overview
If someone truly knows the correct answer, they will remain confident in their ability to answer the question correctly, even when the question is asked in different ways. Inspired by this, we think a model may be overconfident if the model is confident in its ability to answer a question correctly but loses confidence when the question format changes. In such cases, the model’s original confidence should be calibrated. This approach, which reduces overconfidence by leveraging the consistency of the model’s confidence across differently phrased questions, is termed $C^{3}$ (Confidence Consistency-based Calibration). We focus on calibrating the model’s confidence on free-form questions cause they are the most commonly used question format in practice.
@@ 6 Conclusion
In this paper, we first examine LLMs’ ability to assess the factuality of their responses using internal states before and after response generation. Our findings show that LLMs can predict the correctness of their answers prior to generation, providing a cost-efficient approach that avoids inference, with this ability further enhanced post-generation. Next, we introduce $C^{3}$ (Confidence Consistency-based Calibration), a method to refine the model’s perception after response generation. Experimental results demonstrate that $C^{3}$ substantially improves LLMs’ ability to recognize what they do not know and consistently enhances their overall perception.
@@
Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62302486, the Innovation Project of ICT CAS under Grants No. E361140, the CAS Special Research Assistant Funding Project, the project under Grants No. JCKY2022130C039, the Strategic Priority Research Program of the CAS under Grants No. XDB0680102, and the NSFC Grant No. 62441229.
@@ A.5 The Impact of Training Sample Amount
Figure 8: Comparison between Align-A with representative baselines.
A.6 Alignment of Representative Baselines
In this section, we introduce several mainstream confidence estimation methods and compare their alignment scores with those obtained from LLMs’ internal states. The methods are:
  • Probabilistic Confidence (Prob): This method calculates the average generation probability of all tokens in the response as the model’s confidence in the answer (Guo et al., 2017).
  • Verbalized Confidence (Verb) (Xiong et al., 2023): Some studies have shown that LLMs can express their confidence in answers using words.
  • Semantic Consistency-Based Confidence (Consis) (Manakul et al., 2023): Rather than relying on probabilistic confidence, this method measures consistency across multiple responses to the same question, using NLI models or LLMs to check if the responses convey the same meaning.
The Vanilla method in our paper is a strong baseline because prior research (Ni et al., 2024b; Su et al., 2024) has shown that confidence derived from the model’s internal states (the Vanilla method) more accurately reflects the model’s true capabilities compared to other methods like probabilistic confidence, verbalized confidence, and semantic consistency-based confidence. We evaluate these baselines using the following experimental setup:
  • Answer and confidence generation: For Prob and Verb, we obtain the greedy answer. For Consis, we generate the greedy answer and additionally sample 10 responses with a temperature of 0.8. We use Qwen2.5-32B-Instruct (Yang et al., 2024) to evaluate the semantic consistency between each sampled response and the greedy answer.
  • Confidence binarization: We select the optimal threshold using the training (10,000 samples for each dataset) and development sets to binarize the confidence scores.
Figure 8 presents a comparison of the alignment between answer correctness and the confidence estimated by various methods, alongside Align-A. Align-A is trained on the same training set as the baselines (10,000 samples). The results demonstrate that Align-A consistently outperforms all baseline methods across the board. Therefore, in $C^{3}$, we focus exclusively on further improving performance using hidden states.
A.7 Costs of $C^{3}$
Cost of $C^{3}$: 1) Cost of question reformulation: This requires one LLM inference call. In this step, the model generates multiple candidate answers in a single generation, and automated rules are used to construct multi-choice questions from these answers. 2) Cost of consistency checking: This requires $\gamma=4$ (i.e., count of constructed multi-choice questions) LLM inference calls because the model is asked to answer four different multiple-choice questions (i.e., with $k=2,4,6,8$ options). We then obtain the model’s confidence via the MLP and use an automated program to check the consistency of the confidence. Therefore, $C^{3}$ requires only $\gamma+1=5$ LLM inference calls.
Cost of existing semantic consistency-based methods: Semantic consistency-based methods often require generating multiple responses and using powerful LLMs to assess the semantic consistency between texts. In our experiment, generating answers takes 11 inference calls (1 greedy answer + 10 sampling answers), and consistency checking takes 10, totaling 21 inference calls. Compared to existing semantic consistency-based confidence estimation methods that rely on text semantic consistency, $C^{3}$ achieves a much lower cost.